OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (anonymous):

y = mx + b y = 3x + 1 y = -3x + 7

OpenStudy (anonymous):

this system has one solution because it has just one point of interception. If you solve the system you get x = 1 , y = 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

part 2: y = 2x+ 2 y = 2x + 3

OpenStudy (anonymous):

This system has no solutions because they represents parallel lines so they never intersect

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

part 3 : 2y = 2x + 8 y = x + 4

OpenStudy (anonymous):

This system has infinitely many solutions because both equations represent the same line.

OpenStudy (anonymous):

well It was a pleasure to help you Saludos desde Perú

OpenStudy (anonymous):

thank you